# OFFSIDE: Benchmarking Unlearning Misinformation in Multimodal Large Language Models

### v1

@@ 3.3 Evaluation Metrics
OFFSIDE provides a comprehensive evaluation framework for unlearning methods in MLLMs, assessing unlearning efficacy, generalizability, and model utility as defined by (Liu et al., 2024d), along with the model’s ability to integrate with post-training interventions (continual learning). To ensure a comprehensive evaluation, we assess the performance of the vanilla, unlearned, and relearned models on MM-Bench. We only report experimental results for each unlearning method where the model’s general capabilities are not excessively degraded. This approach guarantees that all models maintain their general capabilities throughout the process, allowing for a fair comparison of both forgetting efficacy and functional consistency.
3.3.1 Classification
To test whether the tested model can recall unlearning targets when certain rumors are provided in prompts, we design the classification task using GPT-4o to apply perturbations near the correct answer. Let $a^{n}$ represent the correct answer. Using GPT-4o, we generate a perturbation set $\mathcal{A}^{n}=\{a_{1}^{n},a_{2}^{n},a_{3}^{n},a_{4}^{n}\}$, where the model modifies the correct answer while maintaining its structure and linguistic template. Among these four responses (one correct and three perturbations), only the original is factually correct. Let ${\bm{I}}^{n}$ and ${\bm{Q}}^{n}$ denote the input images and questions, respectively, where $n$ denotes the ID of a given sample. The inputs can be denotes as (${\bm{I}}^{n}$, ${\bm{A}}^{n}$, ${\bm{Q}}^{n}$). The model predicts $\hat{y^{n}}$ by maximizing the probability $P(a^{n}\mid\displaystyle{\bm{I}}^{n},\displaystyle{\bm{A}}^{n},\displaystyle{\bm{Q}}^{n},\theta)$, where $\theta$ is the evaluated model:
$\hat{y^{n}}=\arg\max_{a^{n}_{i}\in{\displaystyle{\bm{A}}^{n}}}P(a^{n}\mid\displaystyle{\bm{I}}^{n},\displaystyle{\bm{A}}^{n},\displaystyle{\bm{Q}}^{n},\theta).$
In the unimodal setting, the input simplifies to $(\emptyset,\displaystyle{\bm{A}}^{n},\displaystyle{\bm{Q}}^{n},\theta)$. To evaluate classification performance, accuracy Acc is computed as following:
$\text{Acc}=\frac{1}{|\displaystyle{\bm{A}}^{n}|}\sum_{a^{n}_{i}\in\displaystyle{\bm{A}}^{n}}\mathbb{I}(\hat{y^{n}}=a^{n}).$
3.3.2 Generation
The generation score used in our paper is defined as the mean of the four evaluation metrics: ROUGE-1, ROUGE-2, ROUGE-L (Lin, 2004), and BLEU (Papineni et al., 2002). Specifically, it is computed as follows:
$\text{Generation Score}=\text{Mean}\left(\text{ROUGE-1}+\text{ROUGE-2}+\text{ROUGE-L}+\text{BLEU}\right).$
3.3.3 Factuality Score
Following previous work (Liu et al., 2024c), we use GPT-4o as an evaluator to assess the factuality, fluency, and semantic relevance of the generated sentences. For each question, we assign a score to the generated answer on a scale from 1 to 10. A score of 1 indicates that the content is completely incorrect or consists of meaningless symbols, while a score of 10 signifies that the answer is factually accurate and well-organized in a coherent sentence.
@@ 4.1 Experiment setup
Training. We employ the Qwen2.5-VL series model as the base model for unlearning. Supervised Fine-Tuning (SFT) is performed using LoRA with a batch size of 4. For methods that require access to the retain set $\mathcal{D}_{\textbf{retain}}$, we adopt a balanced forget-retain update schedule, in contrast to the inner-loop forget and outer-loop retain strategy proposed in (Liu et al., 2024c). Specifically, we use a forget-to-retain step ratio of 1:31\!:\!3 (which corresponds to the size ratio of the forget and retain sets) to enhance training stability during the unlearning process. All experiments are conducted on a single H200 GPU (96GB).
Unlearning Algorithms. We evaluates five representative machine unlearning methods to enable an extensive analysis. Specifically, the methods examined include Gradient Ascent (GA) (Yao et al., 2024a), Gradient Difference (GD) (Liu et al., 2022), KL Minimization (Yao et al., 2024b), Preference Optimization (PO) (Maini et al., 2024), and Negative Preference Optimization (NPO) (Zhang et al., 2024). Since some of these approaches may cause progressive degradation in overall model performance during unlearning, we carefully select and report results only under conditions where the model’s core functionality is preserved, thus ensuring the practical utility of the unlearned model.
4.2 Diverse Experimental Scenarios for OFFSIDE
@@
Selective Unlearning: In this scenario, we focus on removing only the private information of a given image while preserving shared, non-sensitive attributes. Specifically, the shared information of $\mathcal{D}_{\textbf{forget}}$ is removed to $\mathcal{D}_{\textbf{retain}}$ and the left private information serve as the $\mathcal{D}_{\textbf{forget}}$. This approach is more realistic, as it enables the model to maintain its core ability to recognize players based on essential characteristics, such as name, height, and dominant foot.
Corrective Relearning: This setting operates within a continual learning framework, where the unlearned model, $\theta_{u}$, is allowed to relearn the facts. This not only assesses the model utility of $\theta_{u}$ but also evaluates whether the unlearned knowledge can be effectively recovered.
Unimodal Unlearning: In this setup, we combine the name of each entity with ${\bm{Q}}^{n}$. During unlearning, we set the input image to $\emptyset$. This allows us to test whether the LLM unlearning algorithms can seamlessly integrate into multimodal unlearning methods. Additionally, it aids researchers in understanding how MLLMs store knowledge.
Table 2: Results of Complete Unlearning. In this setting, we treat all 14 VQA pairs of an image as rumors. The best results of five baselines are highlighted in blue.
@@
NPO 45.1% 0.371 5.48 49.3% 0.337 3.71 50.2% 0.408 5.69 78.0%
Table 3: Results of Selective Unlearning. In this setting, we merely unlearn the 6 private information of the rumor images and preserve shared ones. The best results of five baselines are highlighted in blue.
@@ 4.3 Experimental Results
Table 2 shows the results of the setting of Complete Unlearning which is a common setting in previous works (Liu et al., 2024c; Dontsov et al., 2024). From this table, GA and NPO results in a significant drop in accuracy on both the test set and retain set while performing the forgetting process. KL and PO demonstrate strong performance on both of the Qwen2.5-VL 7B and 3B models, especially on preventing significant degradation in model performance. However, they are prone to overfitting to the $\mathcal{D}_{\textbf{relearn}}$ in practical. As a result, the methods require very careful control of the training process, limiting their practicality.
Table 3 presents the results of Selective Unlearning. We observe that all the baselines exhibit a performance drop(compared to the vanilla model) in both private information and shared information. This indicates that the tested baselines have trouble selectively unlearning private information in a given image while preserving shared information. This uncovers that existing methods focus on entity-level unlearning, which disrupts all associations between a given image and related text, making it challenging to selectively cut part of the associations.
Table 4 presents the results of Corrective Relearning. The model used here is based on Table 2, where we retrain the unlearned model. Surprisingly, we found that after relearning, all of the baselines exhibit a "bounce-back" effect on either the 3B or 7B model (since relearning serves as a continual learning process, it is common to forget, but uncommon to recover), indicating that the knowledge previously forgotten can be easily recovered through simple retraining. Specifically, KL achieves a fact score of 0.57 on the forget set, which increases to 4.55 after relearning. This suggests that none of the baselines truly forget the rumor information; instead, they merely conceal it.
Figure 3 presents the results of the Unimodal Unlearning setting. In the multimodal setup, the input consists of both text and images, while in the unimodal setup, only text is provided. As shown in the results, multimodal unlearning methods, while forgetting information, simultaneously lead to a trade-off with a decrease in model utility. All unimodal unlearning methods struggle to unlearn multimodal rumors. This suggests that the target information is not only restored in LLMs but also embedded within the visual layer of MLLMs. This highlights the need for researchers to design unlearning methods specifically tailored to the unique characteristics of MLLMs.
@@ 4.4 Discussion
All baselines have trouble unlearning visual rumors. As shown in Figure 4, when confronted with deceptive visual rumors, the model can easily be misled due to its powerful reasoning capabilities. This is intuitive because, even if the model forgets the visual rumors at the visual-text fusion level, it still lacks the knowledge needed to answer the question. Consequently, the model’s response largely depends on the information it perceives in the image, without recognizing that the visual information is unreliable.
@@
All of the tested baselines remain vulnerable to prompt based attacks. Although certain methods achieve low generation and fact scores on the forget set, they still maintain high classification accuracy. This indicates that when rumor information appears in the prompt, the model can still recognize and select the incorrect knowledge, thereby exposing its susceptibility to prompt-induced retrieval. For instance, as shown in Table 2, PO demonstrates strong performance in generation and fact scoring, suggesting effective forgetting. However, its classification accuracy remains close to that of the original, unmodified model, revealing a critical gap in current unlearning approaches. This persistent ability to match forgotten content in classification task underscores the need for more comprehensive and robust unlearning techniques that address both generative and discriminative exposure.
@@
Unlearning efficacy is largely driven by catastrophic forgetting. In Figure 4, we compare the GPT-evaluation results of models relearned after forgetting with those of the directly relearned vanilla model. We observe that the knowledge unlearned by the baselines closely resembles catastrophic forgetting in continual learning. Specifically, the unlearned sample IDs through GA, GD, KL, and NPO show 71%, 48%, 58%, and 60% similarity to the forgotten IDs after a simple relearning step. This suggests that the unlearning ability of the tested baselines is primarily driven by catastrophic forgetting in continual learning. This is largely because these unlearning algorithms can be viewed as a form of continual learning, which inherently results in catastrophic forgetting. This phenomenon demonstrates how catastrophic forgetting can be leveraged as a method for machine unlearning and highlights a promising direction for future research.
@@
Methods such as KL Minimization demonstrate greater effectiveness when applied to a 7B model, but show reduced efficacy with a 3B model. This is primarily due to the random direction of optimization in gradient-ascent-based methods. Before model collapse occurs, these methods struggle to control the optimization direction, which may lead to significant deviations in the results. In contrast, methods like PO, which do not rely on gradient ascent, show more stable performance across both models.
@@ 5 Conclusion
We introduce OFFSIDE, designed to simulate a real-world scenario for unlearning in MLLMs. We propose four distinct settings (Complete Unlearning, Selective Unlearning, Corrective Relearning, and Unimodal Unlearning) to establish a robust unlearning framework and comprehensively evaluate a list of representative machine unlearning baselines. Our findings indicate that: all baselines struggle to unlearn visual rumors, and the unlearned knowledge can be easily recovered through prompt attacks (classification tasks) or simple relearning. Moreover, directly applying unimodal unlearning methods proves inadequate for effectively removing multimodal rumors, highlighting the need for algorithms specifically tailored to MLLM unlearning. Notably, our corrective relearn setting reveals that the unlearning ability of the tested baselines is primarily driven by catastrophic forgetting within continual learning. This suggests a potential connection between machine unlearning and continual learning. Overall, our findings provide valuable empirical insights that guide the development of more effective unlearning methods for future MLLM research.
Ethics statement.
All images used in this study are collected from real-world sources. To prevent the spread of misinformation or rumors, we will provide the original URLs for each image. These URLs are intended solely for research purposes and should be used exclusively by qualified researchers for academic investigation. This transparency ensures traceability and supports reproducibility while upholding responsible research practices.
Reproducibility statement.
The computational resources and experimental setup required for this study are provided. Upon acceptance of the paper, we will release all source code associated with our experiments.
@@
  • Section A: Introduces the baseline methods compared in our experiments.
  • Section B: Introduces the MM-Bench Indicator Definitions.
  • Section C: Describes the supervised Fine-tuning process.
  • Section D: A detailed version of hyperparameters settings.
  • Section E: A case study of our proposed settings.
  • Section F: Outlines the extent to which we employ LLMs.
  • Section G: A detailed description of GPT prompt strategy.
Appendix A Unlearning Methods
Gradient Ascent(GA) Yao et al. (2024b): This method updates the model parameters by maximizing the likelihood of misprediction for the samples in the forget set $D_{\text{forget}}$. For a given sample $x\in D_{\text{forget}}$, the loss function is defined as:
@@ Appendix A Unlearning Methods
Gradient Difference (GD) Liu et al. (2022): This method extends gradient ascent by simultaneously focusing on forgetting the samples in the forget set $D_{\text{forget}}$ while preserving performance on the retain set $D_{\text{retain}}$. The objective is to balance increasing the loss on the forget set and minimizing its impact on the retain set. The overall loss function to be minimized is formulated as:
@@
KL_Min Yao et al. (2024a): This method extends gradient ascent by introducing an additional objective that minimizes the Kullback–Leibler (KL) divergence between the predictions of the original model $M_{\text{ori}}$ and the updated model $M_{\text{new}}$ on the retain set $D_{\text{retain}}$. The KL divergence loss is defined as:
$\mathcal{L}_{\text{KL}}=\frac{1}{|D_{\text{retain}}|}\sum_{s\in D_{\text{retain}}}\frac{1}{|s|}\sum_{i=2}^{|s|}\mathrm{KL}\!\left(M_{\text{ori}}(s_{<i})\;\Big\|\;M_{\text{new}}(s_{<i})\right).$ (3)
@@
Preference Optimization (PO) Maini et al. (2024): This method steers the model to align with newly generated responses such as “I do not know the answer” and its variants for questions belonging to the forget set $D_{\text{forget}}$. At the same time, it incorporates a retain-set term to ensure that predictions on the retain set $D_{\text{retain}}$ remain unaffected. The total objective function is formulated as:
@@
Negative Preference OptimizationZhang et al. (2024): In our work, we adopt the Negative Preference Optimization (NPO) technique to unlearn undesirable data, thereby mitigating the catastrophic collapse often observed in gradient ascent–based methods. NPO builds on the preference optimization framework, but specifically targets negative samples from the forget set $D_{\text{forget}}$.
@@
Appendix B MM-Bench Indicator Definitions

### v2

@@ 3.5 Evaluation Metrics
OFFSIDE provides a comprehensive evaluation framework for unlearning methods in MLLMs, assessing unlearning efficacy, generalizability, and model utility as defined by (liu2024machine), along with the model’s ability to integrate post-training interventions (such as continual learning). We have defined four tasks for evaluation: Classification, Generation, Factuality Score, and MM-Bench tasks. [^4] We have provided a detailed description in Appendix D. Performance on the MM-Bench serves as a disqualifying criterion for selecting experimental results. Experimental results are reported only for those unlearning methods where the model’s general capabilities are not excessively degraded. This ensures that all models maintain their overall functionality throughout the process, allowing for a fair comparison of both forgetting efficacy and functional consistency.
@@ 4.1 Experiment setup
Training. We employ the Qwen2.5-VL series model as the base model for unlearning. Supervised Fine-Tuning (SFT) is performed using LoRA with a batch size of 4. For methods that require access to $\mathcal{D}_{\textbf{retain}}$, we adopt a balanced forget-retain update schedule, in contrast to the inner-loop forget and outer-loop retain strategy proposed in (liu2024protecting). Specifically, we use a forget-to-retain step ratio of 1:31\!:\!3 (which corresponds to the size ratio of the forget and retain sets) to enhance training stability during the unlearning process. All experiments are conducted on a single H20 GPU (96GB).
Unlearning Algorithms. We evaluates five representative machine unlearning methods to enable an extensive analysis. Specifically, the methods examined include Gradient Ascent (GA) (yao2024machine), Gradient Difference (GD) (liu2022continual), KL Minimization (yao2024large), Preference Optimization (PO) (maini2024tofu), Negative Preference Optimization (NPO) (zhang2024negative). Since some of these approaches may cause progressive degradation in overall model performance during unlearning, we carefully select and report results only under conditions where the model’s core functionality is preserved, thus ensuring the practical utility of the unlearned model.
4.2 Experimental Scenarios
@@
Fine-grained Unlearning: In this scenario, we focus on removing only the private information of a given image while preserving shared, non-sensitive attributes. Specifically, the shared information of $\mathcal{D}_{\textbf{forget}}$ is removed to $\mathcal{D}_{\textbf{retain}}$ and the left private information serve as the $\mathcal{D}_{\textbf{forget}}$. This approach is more realistic, as it enables the model to maintain its core ability to recognize players based on essential characteristics, such as name, height, and dominant foot.
Corrective Relearning: This setting operates within a continual learning framework, where the unlearned model, $\theta_{u}$, is allowed to relearn the facts. This not only assesses the model utility of $\theta_{u}$ but also evaluates whether the unlearned knowledge can be effectively recovered. [^5] Previous works merely address the continual unlearning(gao2024large) proble which is quite from our work.
Unimodal Unlearning: In this setup, we combine the name of each entity with questions. During unlearning, we set the input image to empty. This allows us to test whether the LLM unlearning algorithms can seamlessly integrate into multimodal unlearning methods. Additionally, it aids researchers in understanding how MLLMs store knowledge.
@@
NPO 45.1% 0.371 3.22 49.3% 0.337 3.71 50.2% 0.408 5.69 78.0%
Table 2: Results of Complete Unlearning. The best results of five baselines are highlighted in blue.
@@
Table 3: Results of Fine-grained Unlearning. The best results of five baselines are highlighted in blue.
@@ 4.3 Experimental Results
Table 2 shows the results of Complete Unlearning. From this table, GA and NPO results in a significant drop in accuracy on both the test set and retain set while performing the forgetting process. KL and PO demonstrate strong performance on both of the Qwen2.5-VL 7B and 3B models, especially on preventing significant degradation in model performance.
Table 3 presents the results of Fine-grained Unlearning. We observe that all the baselines exhibit a performance drop(compared to the vanilla model) in both private information and shared information. This indicates that the tested baselines have trouble selectively unlearning private information in a given image while preserving shared information. This uncovers that existing methods focus on entity-level unlearning, which disrupts all associations between a given image and related text, making it challenging to be applied to real world applications.
Table 4 presents the results of Corrective Relearning. The model used here is based on Table 2, where we retrain the unlearned model on new data $\mathcal{D}_{\textbf{relearn}}$. Surprisingly, we found that after relearning, all of the baselines exhibit a "bounce-back" effect on either the 3B or 7B model, indicating that the knowledge previously forgotten can be easily recovered through simple retraining. Specifically, KL achieves a fact score of 0.57 on the forget set, which increases to 4.55 after relearning. This suggests that none of the baselines truly forget the rumor information; instead, they merely conceal it. This extends the finding of LLM unlearning(xu2025relearn) to MLLM. [^6] While xu2025relearn utilizes relearning to forget the target, we focus on the rumor recovery after relearning.
Figure 3 presents the results of the Unimodal Unlearning setting. In the multimodal setup, the input consists of both text and images, while in the unimodal setup, only text is provided. As shown in the results, all unimodal unlearning methods struggle to unlearn multimodal rumors. This suggests that the target information is not only restored in LLMs but also embedded within the visual layer of MLLMs. This highlights the need for researchers to design unlearning methods specifically tailored to the unique characteristics of MLLMs.
@@ 4.4 Discussion
All baselines struggle with unlearning visual rumors. We examined all instances of visual rumors and found that none were successfully unlearned by any method. As shown in Figure 4, when faced with deceptive visual rumors, the model is easily misled due to its powerful reasoning capabilities. This is intuitive because, even if the model forgets the visual rumors at the visual-text fusion level, it still lacks the necessary knowledge to correctly answer the question. As a result, the model’s response primarily depends on the information it perceives in the image, without recognizing that the visual information is unreliable. This highlights the need for developing specific algorithm for the visual target.
@@
All of the tested baselines remain vulnerable to prompt based attacks. Although certain methods achieve low generation and fact scores on the forget set, they still maintain high classification accuracy. This indicates that when rumor information appears in the prompt, the model can still recognize and select the incorrect knowledge, thereby exposing its susceptibility to prompt-induced retrieval. For instance, as shown in Table 2, PO demonstrates strong performance in generation and fact scoring, suggesting effective forgetting. However, its classification accuracy remains close to that of the original, unmodified model, revealing a critical gap in current unlearning approaches. This persistent ability to match forgotten content in classification task underscores the need for more robust unlearning techniques.
@@
Unlearning efficacy is largely driven by catastrophic forgetting statistically. In Figure 4, we compare the GPT-evaluation results of models relearned after forgetting with those of the directly relearned vanilla model. We observe that the knowledge unlearned by the baselines closely resembles catastrophic forgetting in continual learning statistically. Specifically, the unlearned sample IDs through GA, GD, KL, and NPO show 71%, 48%, 58%, and 60% similarity to the forgotten IDs after a simple relearning step. This suggests that the unlearning ability of the tested baselines is primarily driven by catastrophic forgetting. This phenomenon demonstrates how catastrophic forgetting can be leveraged as a method for machine unlearning and highlights a promising direction for future research.
@@
Methods such as KL Minimization demonstrate greater effectiveness when applied to a 7B model, but show reduced efficacy with a 3B model. This is primarily due to the random direction of optimization in gradient-ascent-based methods. Before model collapse occurs, these methods struggle to control the optimization direction, which may lead to significant deviations in the results. In contrast, methods like PO, which do not rely on gradient ascent, show more stable performance across both models. [^7] For more discussions, please refer to the Appendix.
5 Limitations
OFFSIDE is the first work to introduce the novel concept of removing visual rumors. However, collecting visual rumors presents a significant challenge, as such rumors are scarce. Specifically, each player is associated with only 8 QA pairs, among which merely one constitutes a visual rumor. Furthermore, while we have identified and discussed several limitations of existing methods, we do not propose a new algorithm capable of effectively addressing these shortcomings. We leave these as promising directions for future research.
@@ 6 Conclusion
We introduce OFFSIDE, designed to simulate diverse real-world scenarios for unlearning in MLLMs. We propose four distinct settings (Complete Unlearning, Fine-grained Unlearning, Corrective Relearning, and Unimodal Unlearning) to establish a robust unlearning framework and comprehensively evaluate a list of representative machine unlearning baselines. Our findings indicate that: all baselines struggle to unlearn visual rumors, and the unlearned knowledge can be easily recovered through prompt attacks (classification tasks) or simple relearning. Moreover, directly applying unimodal unlearning methods fails to remove multimodal rumors. Notably, our corrective relearning setting reveals that the unlearning ability of the tested baselines is primarily driven by catastrophic forgetting. Overall, our findings provide valuable empirical insights that guide the development of more effective unlearning methods for future MLLM MU research.
@@
  • Section A: More details about Related work.
  • Section B: Broader impact of Visual Rumors.
  • Section C: Details of tested baselines.
  • Section D: Details of evaluation metrics.
  • Section E: Introduces the MM-Bench Indicator Definitions.
  • Section E: Vanilla Model Fine-tuning.
  • Section G: Details of experimental settings.
  • Section H: Data construction.
  • Section I: Further findings.
  • Section J: A case study of our proposed settings.
  • Section K: A detailed description of GPT prompt strategy.
  • Section L: Future work.
  • Section M: Discussion.
  • Section N: Use of AI.
Appendix A Extra Related Work
LLM Machine Unlearning. Existing benchmarks in LLM MU have been used to test unlearning in various contexts, such as elimination of personal identification data (patil2023can), copyright protection (eldan2023s) and harmful content removal (lu2022quark). Gradient Ascent (GA) (yao2024large) was introduced to optimize the model parameters so as to maximize the removal of targeted information from the training data. However, GA often degrades performance on the retained set. Subsequent methods, including gradient descent (GD) (liu2022continual), KL-based objectives (yao2024machine; liu2024revisiting), and “I don’t know” (IDK) losses (maini2024tofu), were proposed to exert finer control over the outputs of unlearned models and to mitigate collateral damage. Additionally, Negative Preference Optimization (NPO) (zhang2024negative) reframes LLM unlearning as a preference-optimization problem.
Model Editing. In this subsection, we mainly focus on the difference between Machine Unlearning and Model editing. Model editing aims to update facts in LLMs without costly retraining(decao2021editingfactualknowledgelanguage). Various model editing methods have been proposed, such as ROME (meng2023locatingeditingfactualassociations) and MEMIT (meng2023masseditingmemorytransformer), which show better generalization than naive fine-tuning. Machine unlearning and model editing are two distinct research areas, each with its own data formats and evaluation standards. Model editing focuses on making targeted, precise modifications (preferably with an emphasis on locality) to a model’s behavior or knowledge, while machine unlearning aims to broadly remove specific information, prioritizing overall consistency. Currently, these two areas are typically studied in isolation. Due to their different objectives (e.g., target focus and data types), their evaluation methodologies also differ significantly, despite the fact that the evaluation metrics may be quite similar. Most importantly, both areas are still in the early stages within the context of MLLMs.
Appendix B Social Impacts of Visual Rumors
Impact on Football field. Compared to purely text-based rumors, visual rumors pose additional risks of infringing on an individual’s portrait rights. In addition, misinformation about football transfers of a certain player can have significant real-world consequences. False rumors often lead to emotional reactions from fans, causing unnecessary excitement or disappointment. Unlearning techniques can mitigate these harms by preventing the spread of misinformation and ensuring decision-making is based on verified information.
Generalization. The issue of visual rumors in the football field is not isolated; it can generalize to other domains, such as sports journalism, social media, and financial markets, where rumors are prevalent. Unlearning such rumors is crucial for preserving trust, reducing instability, and promoting more reliable information across various societal sectors. OFFSIDE provides a route for constructing visual rumors for other fields: one can directly inject false text/icon into a singer/politician’s image. Thus forming the visual rumors. In addition, it is easy to collect benign and harmful information of any given people. In this view, the fine-grained unlearning data can be easily collected in other fields. These prove that OFFSIDE is not limited to the football area and can be generalized to any other field because they share the same fundamental logic.
Appendix C Unlearning Methods
Gradient Ascent(GA) yao2024large: This method updates the model parameters by maximizing the likelihood of misprediction for the samples in the forget set $D_{\text{forget}}$. For a given sample $x\in D_{\text{forget}}$, the loss function is defined as:
@@ Appendix C Unlearning Methods
Gradient Difference (GD) liu2022continual: This method extends gradient ascent by simultaneously focusing on forgetting the samples in the forget set $D_{\text{forget}}$ while preserving performance on the retain set $D_{\text{retain}}$. The objective is to balance increasing the loss on the forget set and minimizing its impact on the retain set. The overall loss function to be minimized is formulated as:
@@
KL_Min yao2024machine: This method extends gradient ascent by introducing an additional objective that minimizes the Kullback–Leibler (KL) divergence between the predictions of the original model $M_{\text{ori}}$ and the updated model $M_{\text{new}}$ on the retain set $D_{\text{retain}}$. The KL divergence loss is defined as:
$\mathcal{L}_{\text{KL}}$ $=\frac{1}{|D_{\text{retain}}|}\sum_{s\in D_{\text{retain}}}\frac{1}{|s|}\sum_{i=2}^{|s|}$ (3)
$\quad\mathrm{KL}\!\left(M_{\text{ori}}(s_{<i})\,\Big\|\,M_{\text{new}}(s_{<i})\right).$
@@
Preference Optimization (PO) maini2024tofu: This method steers the model to align with newly generated responses such as “I do not know the answer” and its variants for questions belonging to the forget set $D_{\text{forget}}$. At the same time, it incorporates a retain-set term to ensure that predictions on the retain set $D_{\text{retain}}$ remain unaffected. The total objective function is formulated as:
@@
Negative Preference Optimizationzhang2024negative: In our work, we adopt the Negative Preference Optimization (NPO) technique to unlearn undesirable data, thereby mitigating the catastrophic collapse often observed in gradient ascent–based methods. NPO builds on the preference optimization framework, but specifically targets negative samples from the forget set $D_{\text{forget}}$.
@@
Appendix D Evaluation Metrics
OFFSIDE provides a comprehensive evaluation framework for unlearning methods in MLLMs, assessing unlearning efficacy, generalizability, and model utility as defined by (liu2024machine), along with the model’s ability to integrate with post-training interventions (continual learning). To ensure a comprehensive evaluation, we assess the performance of the vanilla, unlearned, and relearned models on MM-Bench. We only report experimental results for each unlearning method where the model’s general capabilities are not excessively degraded. This approach guarantees that all models maintain their general capabilities throughout the process, allowing for a fair comparison of both forgetting efficacy and functional consistency.
D.1 Classification
To evaluate whether a model can recall unlearning targets when specific rumors are provided in the prompt, we design a multiple-choice classification task with candidates generated by GPT-4o. Let $a^{n}$ denote the ground-truth answer for sample $n$. We construct a candidate set $\mathcal{A}^{n}=\{a^{n}_{0},a^{n}_{1},a^{n}_{2},a^{n}_{3}\}$, where $a^{n}_{0}\equiv a^{n}$ is the correct answer and the remaining three candidates are perturbations that preserve the linguistic template but alter factual content.
Let $\mathbf{I}^{n}$ and $\mathbf{Q}^{n}$ denote the input image and question, respectively. Given $(\mathbf{I}^{n},\mathbf{Q}^{n},\mathcal{A}^{n})$, the evaluated model with parameters $\theta$ predicts
$\hat{y}^{n}=\operatorname*{arg\,max}_{a^{n}_{i}\in\mathcal{A}^{n}}P_{\theta}\!\left(a^{n}_{i}\mid\mathbf{I}^{n},\mathbf{Q}^{n},\mathcal{A}^{n}\right).$ (7)
In the unimodal setting, we remove the image input:
$\hat{y}^{n}=\operatorname*{arg\,max}_{a^{n}_{i}\in\mathcal{A}^{n}}P_{\theta}\!\left(a^{n}_{i}\mid\mathbf{Q}^{n},\mathcal{A}^{n}\right).$ (8)
We report classification accuracy:
$\mathrm{Acc}=\frac{1}{N}\sum_{n=1}^{N}\mathbb{I}\!\left(\hat{y}^{n}=a^{n}\right),$ (9)
where $\mathbb{I}(\cdot)$ is the indicator function.
D.2 Generation
The generation score used in our paper is defined as the mean of the four evaluation metrics: ROUGE-1, ROUGE-2, ROUGE-L (lin-2004-rouge), and BLEU (papineni-etal-2002-bleu). Specifically, it is computed as follows:
Generation Score $=\text{Mean}\Bigl(\text{ROUGE-1}+$ (10)
ROUGE-2 $+\text{ROUGE-L}+\text{BLEU}\Bigr).$
D.3 Factuality Score
Following previous work (liu2024protecting), we use GPT-4o as an evaluator to assess the factuality, fluency, and semantic relevance of the generated sentences. For each question, we assign a score to the generated answer on a scale from 1 to 10. A score of 1 indicates that the content is completely incorrect or consists of meaningless symbols, while a score of 10 signifies that the answer is factually accurate and well-organized in a coherent sentence.
Appendix E MM-Bench Indicator Definitions